A particle of mass 2kg is placed on a rough plane inclined at an angle 35° to the horizontal.
The coefficient of friction is 2
3
. Find the least force parallel to the plane that is required
(i) to hold the particle at rest
(ii) to make the particle slide up the plane.
1
Expert's answer
2020-06-29T18:08:40-0400

"F_{fr}=\mu N"



"\vec{R}=m\vec{g}+\vec{N}+\vec{F_{fr}}+\vec{F}"






(i) According to second Newton's Law


"\vec{R}=0"

"m\vec{g}+\vec{N}+\vec{F_{fr1}}+\vec{F}=0"

"Ox: -\mu N-F+mg\sin\theta=0"

"Oy:N-mg\cos\theta=0"

"F=mg\sin\theta-\mu mg\cos\theta=mg(\sin\theta-\mu \cos\theta)"

"F=2\ kg\cdot9.81\ m/s^2(\sin35\degree-0.23\cos35\degree)\approx7.557 N"

(ii) According to second Newton's Law


"\vec{R}=0"

"m\vec{g}+\vec{N}+\vec{F_{fr2}}+\vec{F}=0"

"Ox: \mu N-F+mg\sin\theta=0"

"Oy:N-mg\cos\theta=0"

"F=mg\sin\theta+\mu mg\cos\theta=mg(\sin\theta+\mu \cos\theta)"

"F=2\ kg\cdot9.81\ m/s^2(\sin35\degree+0.23\cos35\degree)\approx14.950 N"